OpenStudy (phoenixfire):

Factoring a cubic polynomial. \[x^3-108x^2+3564x-34992\] I haven't needed to factor a cubic in a while and have forgotten how. Any help? I know the answers, just can't get them.

OpenStudy (anonymous):

I know how to factor, but explaining it to someone is complicated. I haven't factored in so long either. I'll try to explain it. Sorry if I'm no help.. :/

OpenStudy (maddyx):

http://answers.yahoo.com/question/index?qid=20080417160912AAt8IHC This could help :)

OpenStudy (anonymous):

Sorry I can't help you. >_< I'm a failed Amby on my first day. :c

OpenStudy (anonymous):

\[\begin{array}{l} {x^3} - 108{x^2} + 3564x - 34992 \\ = x({x^2} - 108x + 2916) + 648(x - 54) \\ = x{(x - 54)^2} + 648(x - 54) \\ = (x - 54)({x^2} - 54x + 648) \\ = (x - 54)(x - 36)(x - 18) \\ \end{array}\]

OpenStudy (anonymous):

@drawar from where did you 2916?

OpenStudy (anonymous):

54^2

OpenStudy (anonymous):

I know but it is not found in the original equation.

OpenStudy (anonymous):

Yes but you have to bring it in, and subtract it later, that's the one and only trick when doing factorization of polynomials.

OpenStudy (anonymous):

Where did you subtract it? And in which course did you take this?

OpenStudy (anonymous):

3564=2916+648

OpenStudy (anonymous):

I don't remember which course, I think it's common sense, no?

OpenStudy (anonymous):

Can you please explain the before the last step? I don't get it, how did you reach to that point?

OpenStudy (anonymous):

@drawar

OpenStudy (anonymous):

Use a calculator to get 2 roots